OpenStudy (anonymous):

Part 1: Use the quadratic formula to solve x2 + 7x = –3. (3 points) Part 2: Using complete sentences, explain the process you used. (2 points)

OpenStudy (anonymous):

Part2: explanation - We have to make all equation equal to zero. -As referring to standard form of quadratic equation \[a x^{2} + b x +c = 0\] - Using quadratic formula which is \[\huge {x=\frac{-b \pm \sqrt{b^{2}-4ab}}{2a}}\] to find roots of equation Which is the same as solution of Luis Rivera. regards